OpenStudy (anonymous):

$5,000 is invested in two different accounts yielding 3% and 3.5% interest. The interest earned on the two accounts is $155. How much money was invested at 3% interest? $1,000 $3,000 $4,000 got b

OpenStudy (***[isuru]***):

hey... if the amount invested at 3% is "x" then the amount invested at 3.5% is (5000 -x) so... \[(x \times \frac{ 3 }{ 100 }) + [ (5000 -x) \times \frac{ 35 }{ 1000 }] = 155\]\[30x + 35(5000 - x) = 155000\]\[-5x = 155000 - 175000\]\[x = \frac{ 20000 }{5 }\]\[x = 4000\] hope this will help ya !!